A Year hence a Father will be Five times as Old as his Son. Two years ago , The father was 3 times as old his son will be four years hence. Find their Present age
Found 2 solutions by addingup, MathTherapy:
Answer by addingup(3677)

F + 1 = 5S
F = 5S - 1
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
F - 2 = 3(S + 4)
F - 2 = 3S + 12
F = 3S + 14
~~~~~~~~~~~~~~~~~~~~~~
5S - 1 = 3S + 14
2S = 15
S = 7 1/2 or 7.5
The son is 7 1/2 years old, and the father is:
(5 x 7.5) - 1 = 37.5 - 1 = 36 1/2 years old
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~`
Check
first equality given by the problem:
"A Year hence a Father will be Five times as Old as his Son"
36.5 + 1 = 5(7.5)
37.5 = 37.5 Correct
.
second equality given by the problem:
"Two years ago , The father was 3 times as old his son will be four years hence"
F - 2 = 3(S+4)
36.5 - 2 = 3(7.5 + 4)
34.5 = 3 x 11.5
34.5 = 34.5 Correct

Answer by MathTherapy(10552)


A Year hence a Father will be Five times as Old as his Son. Two years ago , The father was 3 times as old his son will be four years hence. Find their Present age
The other person's answer is WRONG, and is NONSENSICAL!
Don't these people know by now that it's ABNORMAL to have an age that is NOT an INTEGER?
Correct answers: